OpenStudy (anonymous):

Jason has $160. Liam has $120. In how many years will they both have the same amount of money if Jason saves $100 per year and Liam saves $110? Jason $160 $260 $360 $460 $560 Liam $120 $230 $340 $450 $560 Year 0 1 2 3 4 Here is a problem that can also be solved by making a table. Kayla has $50. Nina has $80. In how many months will they both have the same amount of money if Kayla saves $50 per month and Nina saves $40?

OpenStudy (anonymous):

month 0 1 2 3 kayla 50 100 150 200 nina 80 120 160 200

OpenStudy (anonymous):

can you get it from the table ?

OpenStudy (anonymous):

3 years for the first problem I believe

OpenStudy (anonymous):

3months for the Kayla and Nina as well

OpenStudy (texaschic101):

first part.... 160 + 100y = 120 + 110y 160 - 120 = 110y - 100y 40 = 10y 40/10 = y 4 = y first part is 4 years second part.. 50 + 50m = 80 + 40m 50m - 40m = 80 - 50 10m = 30 m = 30/10 m = 3 second part is 3 months

OpenStudy (anonymous):

good job your welcome